# Relative Preference Optimization: Enhancing LLM Alignment through Contrasting Responses across Identical and Diverse Prompts

### v1

@@ 4.1 Experimental Setup
We assessed RPO against a range of alignment methods. These included SFT (Chung et al., 2022) for initial model adaptation, PPO (Schulman et al., 2017) for reinforcement learning fine-tuning, DPO and IPO (Azar et al., 2024) for preference-based model alignment, and KTO (Ethayarajh et al., 2023) as an alternative approach incorporating human value functions. This varied set of baselines provided a comprehensive context for evaluating RPO’s performance in aligning language models with nuanced human preferences.
For these comparisons, we utilized a range of pre-trained large language models, including LLama-2-7/13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023).
We conducted the evaluations of RPO on the validation sets of Anthropic’s Helpful and Harmless (HH) Dataset for dialogue and the OpenAI Summarization Dataset for summarization. To further challenge RPO’s adaptability and general capability in instruction-following, we integrated the AlpacaEval2.0 leaderboard (Li et al., 2023) into our evaluation benchmark. This benchmark comprises a set of 805 diverse and carefully curated prompts, serving as an ideal platform for testing the model’s ability to follow general user instructions accurately and effectively.
@@
Following the experimental settings similar to KTO (Ethayarajh et al., 2023), we maintained a consistent beta value ($\beta=0.1$) throughout all experiments. The training utilized 8 Nvidia A100 GPUs, with a batch size of 64, optimized with RMSProp optimizer (Tieleman and Hinton, 2017). Initially, we trained the SFT, followed by training the subsequent alignment models based on the SFT. For more detailed information on training and evaluation, please refer to Appendix A.
@@ 4.2 Ablation Study
We initiated our investigation with an ablation study aimed at assessing the viability of using semantically related prompts as effective contrastive pairs for preference optimization. Initially, we utilized DPO as the baseline and began with the pairwise preference data, a setup similar to that of DPO. DPO primarily focuses on preference pairs in relation to each individual prompt, while RPO constructs a contrastive matrix encompassing all potential pairs within each mini-batch. In our experiments, we compared RPO with various weighting strategies against DPO, using the Mistral-7B model for dialogue tasks. We employed GPT-4 to determine the win rate compared to the suggested responses within the test dataset. As presented in Table 1, simple reweighting strategies, such as Uniform Weighting (where all contrastive pairs are considered equally important) and Diagonal Weighting (which places more emphasis on diagonal pairwise terms while treating the other non-diagonal terms evenly), yielded slightly worse results than the baseline DPO. This suggests that not all prompt pairs can be deemed effective contrastive pairs for preference optimization. This finding aligns with our intuition, indicating that only semantically related prompt responses can effectively serve as preferred and rejected pairs for each other. Motivated by these observations, we introduced a sentence embedding model to measure the semantic relatedness between prompts. We then utilized this similarity measure to emphasize the preference contrast for highly related pairs. The experiments demonstrated that with Embedding Reweighting, RPO significantly outperformed DPO in both paired and unpaired data scenarios. This success validates the notion that preference contrast between highly related prompts is extremely valuable for aligning human preferences, a process closely mirroring human learning.
@@
Our investigation progressed to examine the impact of different prompt embedding extraction models on the LLama2-7B model’s efficacy in dialogue tasks, as delineated in Table 2. In this comparative study, we scrutinized four distinct models: all-MiniLM-L6-v2 (Wang et al., 2020), known for its efficiency and balance in handling context; sentence-t5-large (Raffel et al., 2020), designed for generating semantically rich embeddings; paraphrase-mpnet-base-v2 (Song et al., 2020), optimized for capturing paraphrastic variations; and all-distilroberta-v1 (Sanh et al., 2019), recognized for its distilled knowledge from larger models. Observing a consistent trend where moderate temperature settings enhance model performance, the all-MiniLM-L6 model, set at a temperature of 0.5, was selected as the benchmark setting for subsequent comparisons against other SoTA models.
@@
Table 1: Ablation study on RPO weighting strategies. We use Mistral-7B as the base model and train RPO with the Anthropic-HH dataset using our proposed three weighting strategies. We employ GPT-4 to determine the win rate compared to the suggested responses within the test dataset. If applied sentence embedding model is set as all-MiniLM-L6-v2.
@@
Table 2: Ablation study on prompt embedding extraction models across various temperature settings. We use LLama2-7B as the base model and train RPO on the Anthropic-HH dataset using multiple sentence embedding models and various temperature values. We employ GPT-4 to determine the win rate compared to the suggested responses within the test dataset.
@@
$\tau$ all-MiniLM-L6-v2 sentence-t5-large paraphrase-mpnet-base-v2 all-distilroberta-v1
0.25 66.80 67.38 62.30 67.58
0.5 68.75 65.23 65.63 67.78
0.75 67.97 65.43 64.45 65.43
1.0 64.06 65.63 66.41 63.67
@@ 4.3 Benchmark Performance
Table 3: win rate on Anthropic-HH and OpenAI Summarization datasets. We conduct a comparative analysis of RPO (unpaired/paired) against state-of-the-art human preference optimization baselines across various datasets, including Anthropic-HH and OpenAI Summarization. We evaluate their performance using metrics such as GPT-4 win rate and AlpacaEval2.0 leaderboard. The evaluation of AlpacaEval2.0 leaderboard is carried out using Mistral-7B, which has been trained on the Anthropic-HH dataset.
@@
Table 3 offers a detailed comparative analysis of win rates for diverse alignment methods applied to the LLama2-7B, LLama2-13B, and Mistral-7B models, addressing tasks across the Anthropic-HH, OpenAI Summarization datasets, and the AlpacaEval2.0 leaderboard. The array of methods evaluated includes SFT, PPO, IPO, DPO, KTO, and our RPO in both unpaired and paired modalities. Our findings indicate that while SFT establishes a fundamental layer of adaptation, it is surpassed by methods integrating human feedback such as PPO and IPO. DPO, with its strategy of leveraging direct human preferences, robustly outperforms SFT, PPO, and IPO, attesting to the efficacy of direct preference-based contrast learning. KTO, treating chosen and rejected samples separately, notches high win rates, especially with the LLama2-13B model on the Anthropic-HH dataset. Yet, it is the RPO approaches that command the highest win rates across the majority of datasets and models, with the RPO-Paired method particularly dominant, exemplified by a 78.52 win rate on the Anthropic-HH dataset with the Mistral-7B model, highlighting the significant benefits of constructing rich contrastive pairs from prompts with semantic similarities.
In light of Mistral-7B’s exceptional performance on dialogue tasks, we solely advanced this model for additional scrutiny on the OpenAI Summarization dataset. The RPO-Unpaired method surprisingly outperformed the paired configuration with win rates of 50.39 and 50.00, respectively, suggesting that the unpaired setting can be remarkably effective, potentially due to the model’s inherent capability to leverage diverse training signals. Additionally, the Mistral-7B model, previously tuned on the Anthropic-HH dataset, was applied to the AlpacaEval2.0 leaderboard to assess its broad generalizability across a spectrum of instruction-following scenarios. Our results on both the Summarization and AlpacaEval benchmarks were highly promising, with the RPO-Unpaired model showcasing a level of performance that not only competes robustly with its paired counterpart but also exemplifies the method’s adaptability. This strong showing emphasizes RPO’s capacity to harness the information content in unpaired data effectively, further confirming the method’s robustness in diverse training scenarios.
@@ 5 Conclusion and Discussion
Limitations & Future Work. RPO’s current limitation lies in its reliance on the quality of the embedding model. To advance, a more generalized approach is needed for constructing contrast pairs in both paired and unpaired preference data scenarios. A promising direction could involve developing algorithms that directly analyze linguistic patterns and contextual similarities within prompts, thus reducing dependence on any specific embedding model and enhancing RPO’s versatility and effectiveness across various data types.
Appendix A Training and Evaluation Details.
@@
The detailed hyperparameters are presented in Table 4; the majority of these parameters are in accordance with the DPO framework (Rafailov et al., 2023). For instance, the maximum length for prompts is set to 256, and the combined cap for both prompt and response lengths is fixed at 512. Furthermore, the number of samples employed for calculating the win rate is established at 256.
@@
Appendix B Algorithm Details
@@
           Calculate contrastive weight: $w_{ij}=\exp\left(-\frac{d_{ij}}{\tau}\right)$
@@
Appendix C The core Python implementation of RPO.

### v2

@@ 4.1 Experimental Setup
We assessed RPO against a range of alignment methods. These included SFT (Chung et al., 2022) for initial model adaptation, PPO (Schulman et al., 2017) for reinforcement learning fine-tuning, DPO and IPO (Azar et al., 2024) for preference-based model alignment, and KTO (Ethayarajh et al., 2024) as an alternative approach incorporating human value functions. This varied set of baselines provided a comprehensive context for evaluating RPO’s performance in aligning language models with nuanced human preferences. For these comparisons, we utilized a range of pre-trained large language models, including LLaMA2-7/13B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023). We conducted the evaluations of RPO on the validation sets of Anthropic’s HH Dataset for dialogue and the OpenAI Summarization Dataset for summarization. To further challenge RPO’s adaptability and general capability in instruction-following, we integrated the AlpacaEval2.0 leaderboard (Li et al., 2023) into our evaluation benchmark. This benchmark comprises a set of 805 diverse and carefully curated prompts, serving as an ideal platform for testing the model’s ability to follow general user instructions accurately and effectively.
@@
Similar to KTO (Ethayarajh et al., 2024), we process preference data in two distinct ways in our experiments. In the paired setting, each batch comprises $N$ triplets $(x,y_{w},y_{l})$. For the unpaired setting, we deconstruct each triplet into two pairs: $(x,y_{w})$ and $(x,y_{l})$, effectively separating the responses into winners and losers. Following a thorough shuffle of the dataset, we extract $N$ instances each of $(x,y_{w})$ and $(x,y_{l})$ to assemble the batch data. It is important to emphasize that the volume of training data utilized in RPO are identical to those in DPO and KTO. In all experiments, both RPO and the baseline consistently utilized a beta value ($\beta=0.1$) and a sampling temperature of 0. The training utilized 8 Nvidia A100 GPUs, with a batch size of 64, optimized with RMSProp optimizer (Tieleman and Hinton, 2017). Initially, we trained the SFT, followed by training the subsequent alignment models based on the SFT. For more detailed information on training and evaluation, please refer to Appendix D.
@@ 4.2 Ablation Study
We initiated our investigation with an ablation study aimed at assessing the viability of using semantically related prompts as effective contrastive pairs for preference optimization. Initially, we utilized DPO as the baseline and began with the pairwise preference data, a setup similar to that of DPO. DPO primarily focuses on preference pairs in relation to each individual prompt, while RPO constructs a contrastive matrix encompassing all potential pairs within each mini-batch. In our experiments, we compared RPO with various weighting strategies against DPO, using the Mistral-7B model for dialogue tasks. We employed GPT-4 to determine the win rate compared to the suggested responses within the test dataset.
As presented in Table 1, simple reweighting strategies, such as Uniform Weighting (where all contrastive pairs are considered equally important) and Diagonal Weighting (which places more emphasis on diagonal pairwise terms while treating the other non-diagonal terms evenly), yielded slightly worse results than the baseline DPO. This suggests that not all prompt pairs can be deemed effective contrastive pairs for preference optimization. This finding aligns with our intuition, indicating that only semantically related prompt responses can effectively serve as preferred and rejected pairs for each other. Motivated by these observations, we introduced a sentence embedding model to measure the semantic relatedness between prompts. We then utilized this similarity measure to emphasize the preference contrast for highly related pairs. The experiments demonstrated that with Embedding Reweighting, RPO significantly outperformed DPO in both paired and unpaired data scenarios. This success validates the notion that preference contrast between highly related prompts is extremely valuable for aligning human preferences, a process closely mirroring human learning.
@@
Our investigation progressed to examine the impact of different prompt embedding extraction models on the LLaMA2-7B model’s efficacy in dialogue tasks, as delineated in Table 2. In this comparative study, we compared three distinct models: all-MiniLM-L6-v2 (Wang et al., 2020), known for its efficiency and balance in handling context; sentence-t5-large (Raffel et al., 2020), designed for generating semantically rich embeddings; and all-distilroberta-v1 (Sanh et al., 2019), recognized for its distilled knowledge from larger models. Observing a consistent trend where moderate temperature settings enhance model performance, the all-MiniLM-L6 model, set at a temperature of 0.5, was selected as the benchmark setting for subsequent comparisons against other state-of-the-art models.
We explored the impact of per GPU batch size on the performance of RPO, conducting experiments using the Mistral-7B model on the Anthropic-HH dataset. The results of these experiments are presented in Table 3. Note the baseline DPO model, judged by GPT-4, achieved a win-rate of 72.26. When operating with the minimal batch size of 2 per GPU, RPO demonstrates a slight underperformance in comparison to the DPO baseline, which is trained with a batch size of 8 per GPU. This observed discrepancy can largely be ascribed to the reduced dimensions of the similarity matrix. However, as the batch size increases, RPO’s performance consistently improves, surpassing the DPO baseline comfortably at batch sizes 4 and above. This trend highlights the effectiveness of RPO’s approach in leveraging larger amounts of comparative data for preference learning. It demonstrates that while RPO can work with smaller batches, its strength is amplified when provided with a richer and more varied similarity matrix, which is intrinsic to larger batch sizes. For further details on ablation studies concerning the inputs used for computing relative weights, as well as experiments exploring the effects of different beta values and sampling temperatures, please refer to Appendix F.
@@
Table 1: Ablation study on RPO weighting strategies. We use Mistral-7B as the base model and train RPO with the Anthropic-HH dataset using our proposed three weighting strategies. If applied sentence embedding model is set as all-MiniLM-L6-v2.
@@
Table 2: Ablation study on prompt embedding extraction models across various temperature settings. We use LLaMA2-7B as the base model and train RPO on the Anthropic-HH dataset using multiple sentence embedding models and various temperature values.
@@
$\tau$ all-MiniLM-L6-v2 sentence-t5-large all-distilroberta-v1
0.25 66.80 67.38 67.58
0.5 68.75 65.23 67.78
0.75 67.97 65.43 65.43
Table 3: Batch-size ablation Study for RPO on the Anthropic-HH dataset with Mistral-7B model.
Per GPU Batch Size 2 4 6 8
RPO Win Rate (%) 71.48 74.60 74.61 78.52
@@ 4.3 Benchmark Performance
Table 4: Win rate on Anthropic-HH and OpenAI Summarization datasets. We conduct a comparative analysis of RPO (unpaired/paired) against state-of-the-art human preference optimization baselines. We evaluate their performance using metrics such as GPT-4 win rate and AlpacaEval2.0 leaderboard. The evaluation of AlpacaEval2.0 leaderboard is carried out using Mistral-7B, which has been trained on the Anthropic-HH dataset.
@@
Table 4 offers a detailed comparative analysis of the win rates for diverse alignment methods applied to the LLaMA2-7B, LLaMA2-13B, and Mistral-7B models, addressing tasks across the Anthropic-HH, OpenAI Summarization datasets, and the AlpacaEval2.0 leaderboard. The array of methods evaluated includes SFT, PPO, IPO, DPO, KTO, and our RPO in both unpaired and paired modalities. Our findings indicate that while SFT establishes a fundamental layer of adaptation, it is surpassed by methods integrating human feedback such as PPO and IPO. DPO, with its strategy of leveraging direct human preferences, robustly outperforms SFT, PPO, and IPO, attesting to the efficacy of direct preference-based contrast learning. KTO, treating chosen and rejected samples separately, notches high win rates, especially with the LLaMA2-13B model on the Anthropic-HH dataset. Yet, it is the RPO approaches that command the highest win rates across the majority of datasets and models, with the RPO-Paired method particularly dominant, exemplified by a 78.52 win rate on the Anthropic-HH dataset with the Mistral-7B model, highlighting the significant benefits of constructing rich contrastive pairs from prompts with semantic similarities.
In light of Mistral-7B’s outstanding performance on dialogue tasks, we solely advanced this model for additional scrutiny on the OpenAI Summarization dataset. The RPO-Unpaired method surprisingly outperformed the paired configuration with win rates of 50.39 and 50.00, respectively, suggesting that the unpaired setting can be remarkably effective, potentially due to the model’s inherent capability to leverage diverse training signals. Additionally, the Mistral-7B model, previously tuned on the Anthropic-HH dataset, was applied to the AlpacaEval2.0 leaderboard to assess its broad generalizability across a spectrum of instruction-following scenarios. Our results on both the Summarization and AlpacaEval benchmarks were highly promising, with the RPO-Unpaired model showcasing a level of performance that not only competes robustly with its paired counterpart but also exemplifies the method’s adaptability. This strong showing emphasizes RPO’s capacity to harness the information content in unpaired data effectively, further confirming the method’s robustness in diverse training scenarios.
@@ 5 Conclusion and Discussion
Limitations & Future Work. RPO currently faces three limitations: First, it depends on the quality of the embedding model used to construct contrast pairs. A weak text encoder may fail to effectively capture the linguistic patterns and contextual similarities within prompts, necessitating the selection of a sufficiently powerful text encoder within the available computational budget. Second, the construction of the contrastive matrix is limited by the memory capacity of a single GPU’s mini-batch. Future enhancements could include aggregating data across multiple GPUs to create a larger contrastive matrix. Third, the algorithm currently assumes a constant $Z(x)$ for all $x$. Advancing RPO could involve dynamically modeling $Z(x)$ for different prompts, potentially using a neural network, to more effectively handle variations in data inputs.
Appendix A Discussion of $Z(x)$.
Similar to DPO [Rafailov et al., 2023], we define $Z(x)$ as:
$Z(x)=\sum_{y}[\pi_{ref}(y|x)\exp(\frac{1}{\beta}\cdot r(x,y))].$ (15)
However, under the frameworks of Plackett-Luce [Plackett, 1975, Luce, 2005] and Bradley-Terry [Bradley and Terry, 1952], the normalization term $Z(x)$ in DPO is treated as a nuanced constant and can be omitted [Rafailov et al., 2023]. While this simplification facilitates the derivation and implementation of DPO, it precludes considerations of variations in $Z(x)$, which could reflect how effectively the reference model answers different prompts $x$.
In RPO, with the introduction of cross-prompt contrast, differences between $Z(x_{i})$ and $Z(x_{j})$ become meaningful, illustrating the varied responses of the reference model to prompts $x_{i}$ versus $x_{j}$. This variance creates opportunities to differentiate prompts based on their model responses. Estimating $Z(x)$ accurately, however, remains challenging. Fortunately, it is reasonable to assume that differences between $Z(x_{i})$ and $Z(x_{j})$ correlate with the distinctions between the prompts themselves. The weighting strategy introduced in Section 3.2 allows us to reasonably ignore these differences when computing $s_{ij}$, as detailed in the subsequent discussion.
Firstly, the assumption that $Z(x_{i})\approx Z(x_{j})$ is reasonable when $x_{i}$ and $x_{j}$ are similar. Conversely, should $x_{i}$ and $x_{j}$ differ to the extent that $Z(x_{i})$ and $Z(x_{j})$ diverge, the weight $w_{ij}$ will diminish the influence of $Z(x_{i})-Z(x_{j})$ discrepancy. Elaborating further, in instances where $Z(x_{i})\neq Z(x_{j})$ within RPO, $s_{ij}$ as defined in Eq. 12 must be adjusted to:
$\tilde{s}_{ij}=\beta w_{ij}\left(\log\frac{\pi_{\theta}(y_{w,i}|x_{i})e^{Z(x_{i})}}{\pi_{ref}(y_{w,i}|x_{i})}-\log\frac{\pi_{\theta}(y_{l,j}|x_{j})e^{Z(x_{j})}}{\pi_{ref}(y_{l,j}|x_{j})}\right),$ (16)
resulting in a modified loss expression:
$\tilde{L}_{RPO}=-\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\log\sigma(\tilde{s}_{ij})=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\log(1+\exp(-s_{ij}-\beta w_{ij}(Z(x_{i})-Z(x_{j}))).$ (17)
This revised formulation accounts for potential disparities in $Z(x)$ values across different prompts while mitigating their impact through appropriate weighting.
Comparing $\log(\sigma(\tilde{s}_{ij}))$ and $\log(\sigma(s_{ij}))$ elucidates the impact of considering $Z(x)$ in the loss formulation:
For $\log(\sigma(s_{ij}))$, we have:
$\log(\sigma(s_{ij}))=\log\frac{\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}\right)^{\beta w_{ij}}}{1+\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}\right)^{\beta w_{ij}}}$ (18)
For $\log(\sigma(\tilde{s}_{ij}))$, it becomes:
$\log(\sigma(\tilde{s}_{ij}))=\log\frac{\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}\right)^{\beta w_{ij}}e^{\beta(Z(x_{i})-Z(x_{j}))w_{ij}}}{1+\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}\right)^{\beta w_{ij}}e^{\beta(Z(x_{i})-Z(x_{j}))w_{ij}}}$ (19)
Simplified, this can be represented as:
$\log\frac{\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}e^{Z(x_{i})-Z(x_{j})}\right)^{\beta w_{ij}}}{1+\left(\frac{\frac{\pi_{\theta}(y_{w,i}|x_{i})}{\pi_{ref}(y_{w,i}|x_{i})}}{\frac{\pi_{\theta}(y_{l,j}|x_{j})}{\pi_{ref}(y_{l,j}|x_{j})}}e^{Z(x_{i})-Z(x_{j})}\right)^{\beta w_{ij}}}$ (20)
When $Z(x_{i})<Z(x_{j})$, indicating that $\pi_{ref}$ performs better in response to $x_{j}$ compared to $x_{i}$, Eq. 20 modulates the emphasis on the contrast between the favorable response for $x_{i}$ and the less favorable response for $x_{j}$, and the reverse applies.
Eq. 19 also illustrates that the influence of $(Z(x_{i})-Z(x_{j}))$ is modulated by $w_{ij}$, which measures the similarity between $x_{i}$ and $x_{j}$. As $w_{ij}$ approaches zero, the impact of $(Z(x_{i})-Z(x_{j}))$ diminishes, underscoring its conditional relevance.
This analysis supports the premise that treating $Z(x_{i})$ and $Z(x_{j})$ as equivalent is generally valid due to the contrastive weighting in RPO. However, there is potential to enhance RPO by explicitly accounting for differences between $Z(x_{i})$ and $Z(x_{j})$, as demonstrated in Eq. 20. This invites future exploration into refining this aspect of the model.
Appendix B Impact Statements
Relative Preference Optimization (RPO) enhances large language models (LLMs) by aligning them more closely with diverse human preferences. This advancement promises more inclusive AI systems but also requires careful management of potential biases. Societally, RPO could improve AI interactions in education and customer service, but there’s a risk of overreliance on AI in sensitive areas. Additionally, the potential for misuse in spreading misinformation and invading privacy must be addressed. Ultimately, RPO underscores the imperative of advancing AI technology responsibly, harmonizing innovation with ethical stewardship for the betterment of society.
Appendix C Utilization of LLM in Manuscript Preparation
We employed GPT-4 to refine the grammar and structure of our manuscript. This enhanced the readability and organization of the paper.
Appendix D Training and Evaluation Details.
@@ Appendix C Utilization of LLM in Manuscript Preparation
The detailed hyperparameters are presented in Table 5; the majority of these parameters are in accordance with the DPO framework [Rafailov et al., 2023]. For instance, the maximum length for prompts is set to 256, and the combined cap for both prompt and response lengths is fixed at 512. Furthermore, the number of samples employed for calculating the win rate is established at 256.
@@
Appendix E Algorithm Details
@@
           Calculate contrastive weight: $w_{ij}=\text{softmax}\left(-\frac{d_{ij}}{\tau}\right)$
@@
Appendix F Ablation Study
In all ablation studies, we emphasize that sampling was conducted on the Anthropic-HH test dataset using 256 samples with Mistral-7B, and win rates were computed using GPT-4.
F.1 Ablation study on similarity weighting: prompt-only vs. integrated prompt-response
Initially, our model computed weights based solely on prompt similarities. To explore the potential benefits of a more nuanced similarity assessment, we also developed a method that integrates both prompts and responses as a single unit for this computation. The comparative results on the Anthropic-HH test dataset and the AlpacaEval 2.0 benchmark are detailed in Table 6. This analysis reveals that while integrating prompt and response information can enhance performance on specific datasets, it may also reduce the model’s generalizability across broader benchmarks. Given the critical importance of maintaining a balance between detailed embeddings and broad applicability, our final strategy prioritizes using prompt-based similarity computations for weight calculation.
Table 6: Comparative win rates for prompt-only vs. integrated prompt-response weighting strategies.
Method Dataset
Anthropic-HH AlpacaEval 2.0
Prompt-Only 78.52 38.88
Integrated Prompt-Response 80.86 35.47
F.2 Ablation study on $\beta$ Values
In our experiments, a beta value of 0.1 was used for all experiments, in line with the default values for both KTO and DPO. We subsequently explored alternative beta values of 0.5 and 0.8 to assess their impact. Note for the best performance of RPO, the $\tau$ should be adjusted accordingly when $\beta$ changes. Here, we keep the same $\tau=0.5$. In Table 7, we observe that RPO consistently outperforms the other methods for all $\beta$ values.
Table 7: Performance comparison of DPO, KTO, and RPO at different beta values.
Method Beta
0.1 0.5 0.8
DPO 72.26 64.85 58.59
KTO 61.13 65.23 63.28
RPO 78.52 68.36 63.67
F.3 Ablation study on sampling temperature
All experiments were conducted with a default sampling temperature of 0, as detailed in Table 5. This setting aligns with those used in the DPO and KTO methods. Furthermore, we conducted a temperature ablation study, the results of which are presented in Table 8. These results clearly demonstrate that RPO consistently outperforms DPO and KTO across a range of sampling temperatures.
Table 8: Performance comparison of DPO, KTO, and RPO under various sampling temperatures.
Method Temperature
0.0 0.3 0.7
DPO 72.26 70.70 72.65
KTO 61.13 57.81 60.94
RPO 78.52 72.48 75.39
Appendix G The core Python implementation of RPO.